myckboard Le x | 1 Francial State × Fnancial Acco xren How to Course: ACCT × Financial Acco Χ ti-integration-BBLEARN/lapp/content/launchHandler?course ids 132178 1&content id _50741471 us Powered by B eCampus ePortfolios Community Content Collection ents > HW Assignment 1 Menu 5 Answer saved Marked out of 3.00 P Flag question Using the Basic Accounting Equation Henderson Company had beginning-of-the-year total assets of $300,000 and total liabilities of $180,000. a. If during the year total assets increased by $15,000 and total liabilities increased by $40,000, what is the end-of-year total stockholders equity? Fin b. If during the year total assets increased by $60,000 and total liabilities decreased by $5,000, what is the end-of-year total stockholders equity? c. If during the year total liabilities increased by $40,000 and total stockholders equity increased by $35,000, what are the end-of-year total assets? Previous page Next page MacBook Air 3
Answer #1

Answer a:

Stockholders' Equity = Total assets - Total Liability

Beginning of the year Stockholders' Equity = $300,000 - $180,000 = $120,000

End of the year Stockholders' Equity = ($300,000 + $15,000) - ($180,000 + $40,000) = $95,000

End of the year Stockholders Equity $95,000

Answer b:

End of the year Stockholders' Equity = ($300,000 + $60,000) - ($180,000 - $5,000) = $185,000

End of the year Stockholders Equity $185,000

Answer c:

Total Assets = Total liabilities + Stockholders' equity

Beginning of the year equation:

=> $300,000 = $180,000 +$120,000

End of the year Total assets = ($180,000 + $40,000) + ($120,000 +$35,000) = $375,000

End of the year Total assets $375,000